OpenStudy (anonymous):

given that f(x) =(∛x-1)/(x-1) then its limit as x approaches 1 is equal to? a.0 b.1 c.-1 d.does not exist

OpenStudy (anonymous):

is (x-1) all under the cube root?

OpenStudy (anonymous):

No option seem correct

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x-1 not under the cube root

myininaya (myininaya):

\[\text{ let } u=x^\frac{1}{3} ; x->1 => u->(1)^\frac{1}{3}=1\] so we have \[\lim_{u \rightarrow 1}\frac{u-1}{u^3-1}=\lim_{u \rightarrow 1}\frac{u-1}{(u-1)(u^2+u+1)}\]

myininaya (myininaya):

that should help out

OpenStudy (nenadmatematika):

solution is 1/3 if you wrote it correctly

OpenStudy (anonymous):

so if theres no answer i put E.

OpenStudy (nenadmatematika):

I suggest you type it again more carefuly and post again...so I would be sure

OpenStudy (anonymous):

It's 0, look at myininaya's solution... \[\frac{1-1}{1^2+1+1}=\frac{0}{3}=0\]

myininaya (myininaya):

hey rick that top part cancels with (x-1) though

myininaya (myininaya):

i mean u-1

OpenStudy (anonymous):

ah right, my mistake

myininaya (myininaya):

so you do get 1/3

OpenStudy (nenadmatematika):

Nenad is the champ :D

OpenStudy (nenadmatematika):

hahaha...kidding

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this one has a sqr (∛x-1) the other 1 hast no cube root (x-1) and i feel nenad is the champ XD + myiniaya XD

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{\sqrt[3]{x-1}}{x-1}\] are you saying this is the problem?

OpenStudy (anonymous):

that's what i asked him earlier..he said no

OpenStudy (anonymous):

YES THATS the one thanks

OpenStudy (anonymous):

sorry rick im confused

OpenStudy (anonymous):

so what is my answer d or e? e for error

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{\sqrt[3]{x-1} \cdot (x-1)^3}{(x-1) \cdot (x-1)^3}\] \[\lim_{x \rightarrow 1}\frac{(x-1)}{(x-1)(x-1)^3}=\lim_{x \rightarrow 1}\frac{1}{(x-1)^4}\] I get the limit does not exist

OpenStudy (nenadmatematika):

does not exist

OpenStudy (anonymous):

or \[\infty\]

OpenStudy (anonymous):

thanks rick

OpenStudy (anonymous):

d is wrong to i got the result of my question i got 16/20 =90 ^^ thanks to all of u